What will be the wavenumber of yellow radiation having wavelength 240 nm

Grade:11

1 Answers

Arun
5 years ago
Wavelength of light= 2400x10^-10m
frequency is given by=c/wavelength
3x10^8/2400x10^-10
3x10^8/24x10^-8
0.125x10^16
wave number is reciprocal of wavelength
 wave number=1/wavelength=1/2400x10^-10=10^10/2400=0.0416x10^8 m^-1